# Replica Thermodynamic Trade-off Relations: Entropic Bounds on Network Diffusion and Trajectory Observables

### v1

@@ Abstract
We introduce replica Markov processes to derive thermodynamic trade-off relations for nonlinear functions of probability distributions. In conventional thermodynamic trade-off relations, the quantities of interest are linear in the underlying probability distribution. Some important information-theoretic quantities, such as Rényi entropies, are nonlinear; however, such nonlinearities are generally more difficult to handle. Inspired by replica techniques used in quantum information and spin-glass theory, we construct Markovian dynamics of identical replicas and derive a lower bound on relative moments in terms of the dynamical activity. We apply our general result to two scenarios. First, for a random walker on a network, we derive an upper bound on the Rényi entropy of the position distribution of the walker, which quantifies the extent of diffusion on the network. Remarkably, the bound is expressed solely in terms of escape rate from the initial node, and thus depends only on local information. Second, we consider trajectory observables in Markov processes and obtain an upper bound on the Rényi entropy of the distribution of these observables, again in terms of the dynamical activity. This provides an entropic characterization of uncertainty that generalizes existing variance-based thermodynamic uncertainty relations.
Introduction
@@
In quantum information and statistical mechanics, one often needs to compute expectation values of nonlinear physical quantities. For example, purity requires the square of the density operator, $\mathrm{Tr}[\rho^{2}]$. Although purity serves as a key quantity for distinguishing quantum from classical behavior, its experimental measurement is known to be difficult. The swap trick [1] is a technique to measure purity without performing state tomography. The swap trick uses the following identity:
@@
where $\mathrm{SWAP}$ denotes the swap operator. The identity given in Eq. (1) shows that the purity $\mathrm{Tr}[\rho^{2}]$ can be obtained by preparing two replicas of $\rho$ and measuring the expectation value of the SWAP operator acting on $\rho\otimes\rho$. This avoids full state tomography of $\rho$, which would require many measurement shots to reliably estimate $\rho^{2}$. The use of replicas has been highly effective in classical shadow tomography [2, 3, 4] or in the study of the barren plateau in parametrized quantum circuits [5]. The application of hypothetical replicas to the evaluation of nonlinear quantities has a long-standing history. In spin glasses [6, 7, 8], for example, we often need to evaluate the disorder average of the logarithm of the partition function $\mathbb{E}[\ln Z]$, which is difficult to compute directly. The replica method addresses this by introducing $n$ replicas of the system and expressing the logarithm as a limit $\ln Z=\lim_{n\to 0}(Z^{n}-1)/n$. This technique has been widely used in the analysis of learning theory [9, 10, 11, 12, 13] and conformal field theory [14, 15] . These examples illustrate the power of using replicas to calculate expectations of nonlinear quantities.
In this paper, we consider a Markov process consisting of identical replica processes in the context of thermodynamic trade-off relations. Over the past decade, thermodynamic trade-off relations have been extensively studied in stochastic and quantum thermodynamics. In the thermodynamic uncertainty relations [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], trajectory observables are considered and lower bounds on their relative variance are derived, where the bounds involve the entropy production and the dynamical activity. Extensions of these uncertainty relations to the quantum domain have been actively studied [32, 33, 36, 37, 38, 39]. Related work has also examined how to compute state-dependent expectation values [40, 41]. What these relations have in common is that the quantities being computed are linear in the underlying probabilities. For example, in thermodynamic uncertainty relations, we are often interested in quantities of the form
$\mathbb{E}[f]=\sum_{\Gamma}\mathcal{P}(\Gamma)\,f(\Gamma),$ (2)
where $\Gamma$ denotes a stochastic trajectory, $\mathcal{P}(\Gamma)$ is the trajectory probability, and $f(\Gamma)$ is an observable like a current. Equation (2) is linear with respect to probability $\mathcal{P}(\Gamma)$. Typical examples of Eq. (2) include dissipated heat and stochastic displacement. In contrast, nonlinear functions of the probabilities are generally much harder to compute, although they play a crucial role. A prominent example is the Rényi entropy, which measures the degree of uncertainty of a random variable. The thermodynamic uncertainty relation expresses the uncertainty using the variance of an observable. In information theory, however, uncertainty is often quantified using entropy. When the output is categorical rather than numerical, the variance is not a natural measure of uncertainty. Entropy, by contrast, applies directly to discrete outcomes and can be used to quantify uncertainty in this setting. For example, when quantifying the extent that a random walker on a network has propagated over the network, variance is not suitable because the states are labels rather than numerical values.
Specifically, we derive a lower bound on relative moments in terms of the dynamical activity [Eq. (12)], which is referred to as replica trade-off relation. This relation holds only in the replica process. However, by considering specific scenarios, the replicas become virtual, as in the previously mentioned spin-glass case, and we can derive a trade-off relation that holds in an original single process. Specifically, we consider a random walker on a network whose dynamics obey a continuous-time Markov process. Using the replica trade-off relation, we derive trade-off relations in the Rényi entropy for the probability distribution of the walker position, which quantifies the extent of the diffusion of the walker. The derived bound indicates that the Rényi entropy is bounded from above by the dynamical activity at the initial time, which corresponds to the local escape rate from the initial node and can be computed as the sum of the transition rates from the initial node to all nodes reachable in a single jump. In particular, no matter how large the overall network is, it can be computed using only the local information around the initial node. In the second scenario, we consider trajectory observables in a continuous-time Markov process and consider the probability distribution of the trajectory observables. Using the replica trade-off relation, we obtain the upper bound on the Rényi entropy of the observable distribution, where the upper bound involves the dynamical activity, which quantifies the uncertainty of the observable trajectory. The uncertainty in the trajectory observable has been conventionally studied using the variance, and it was shown that its relative variance is bounded from below by the reciprocal of the dynamical activity. Our result shows that not only the relative variance but also uncertainty based on Rényi entropy is fundamentally constrained by the dynamical activity.
@@
Figure 1: (a) Single process scenario. This scenario is usually considered in stochastic thermodynamics. Given a Markov process, we consider stochastic trajectories, which are a random realization of the Markov process. We are interested in stochastic quantities associated with trajectories, whose statistics are linear in the trajectory probability. (b) $K=2$ replica process scenario. We consider two identical Markov processes. The state space is a product space of each Markov process and the trajectories are defined on the product space. We are interested in stochastic quantities of the replica process, which become nonlinear quantities in the original single process.
Results
Replica Markov processes
@@ Results
where the diagonal elements are defined by $W_{\mu\mu}=-\sum_{\mu^{\prime}\neq\mu}W_{\mu^{\prime}\mu}$. In stochastic thermodynamics, the trajectories generated by the master equation [Eq. (3)] are fundamental for analyzing the system’s stochastic dynamics. Several quantities, such as state displacement, dissipated heat, and Shannon entropy, are considered at the level of the stochastic trajectories. We are interested in the dynamics within the time interval $0,\tau]$ ($\tau>0$). Suppose that there are $J$ jump events within the interval. Let $t_{j}$ be the time of the $j$-th jump and let $X_{j}\in\mathfrak{B}$ be the state after the $j$-th jump. The trajectory $\Gamma$ is then expressed by
@@
where $t_{0}=0$ is the initial time and $X_{0}$ is the initial state.
As mentioned in the introduction, we consider replicas of the original process and regard the whole Markov process as a combined process (Fig. 1(b)). For simplicity, here we introduce the case of two replicas. That is, there are two identical processes, where there is no interaction between them. It is straightforward to extend this case to scenarios where there are $K$ replica processes. The state space is a product of the two state sets. The set of states in the combined Markov process of two replicas is given by $\mathfrak{B}^{(\otimes 2)}\equiv\{(B_{1}^{(1)},B_{1}^{(2)}),(B_{1}^{(1)},B_{2}^{(2)}),\ldots,(B_{D}^{(1)},B_{D}^{(2)})\}$. Therefore, the number of total states in the combined process is $|\mathfrak{B}^{(\otimes 2)}|=D^{2}$. From now on, when a superscript $(k)$ is used, it denotes a quantity associated with the $k$-th process, and when a superscript $(\otimes K)$ is used, it denotes a quantity for the combined system in which all the processes $1,2,\dots,K$ are regarded as a single combined process. When we omit the superscript, the corresponding variables refer to those in the single process (i.e., $K=1$ case).
Let $p_{\mu\nu}^{(\otimes 2)}(t)$ be the probability that the combined system is in state $(B_{\mu}^{(1)},B_{\nu}^{(2)})\in\mathfrak{B}^{(\otimes 2)}$ at time $t$. Let $W_{\nu\mu}^{(k)}$ be the transition rate from state $B_{\mu}^{(k)}$ to state $B_{\nu}^{(k)}$ in the $k$-th process. Since we are considering identical processes, $W_{\mu^{\prime}\mu}^{(k)}=W_{\mu^{\prime}\mu}$, but we keep the superscript for clarity. The master equation corresponding to the combined process of the two replicas is given by
@@
Replica trade-off relations
Introducing the combined process with two replicas, we will focus on general scenario in which there are $K$ replicas. We now introduce observables associated with trajectories. Let $N^{(\otimes K)}(\Gamma^{(\otimes K)})$ be an observable of the trajectory $\Gamma^{(\otimes K)}$, which is the trajectory in the combined process of $K$ replicas. Let $\Gamma_{\varnothing}^{(\otimes K)}$ be arbitrary trajectory in the combined process in which there is no jump. For example, $\Gamma_{\varnothing}^{(\otimes K)}$ could be
$\Gamma_{\varnothing}^{(\otimes 2)}=\bigl[(t_{0},(B_{1}^{(1)},B_{1}^{(2)}))\bigr],$ (10)
which consists only of the initial state $(B_{1}^{(1)},B_{1}^{(2)})$. The observable $N^{(\otimes K)}(\Gamma_{\varnothing}^{(\otimes K)})$ can be arbitrary, provided it satisfies
$N^{(\otimes K)}(\Gamma_{\varnothing}^{(\otimes K)})=0.$ (11)
A simple example of such an observable in the single process is a counting observable, that count the number of jumps within the time interval $0,\tau]$. Apparently, such an observable is zero for $\Gamma_{\varnothing}$ and thus it satisfies the condition of Eq. (11).
Let $\mathbb{E}[X]$ denote the expectation value of a random variable $X$. Since $\Gamma$ is a random variable, observable $N^{(\otimes K)}(\Gamma^{(\otimes K)})$ is a random variable as well. We then obtain our first result (see Methods for the derivation):
$\frac{\mathbb{E}\left[|N^{(\otimes K)}|^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[|N^{(\otimes K)}|^{r}\right]^{s/(s-r)}}\geq\frac{1}{1-e^{-\tau K\mathfrak{a}(t=0)}},$ (12)
where $0<r<s$ and $\mathfrak{a}(t=0)$ is the dynamical activity:
@@
Note that the dynamical activity $\mathfrak{a}(t)$ in Eq. (13) is a quantity of a single process [Eq. (3)]; the definition is the same as conventional Markov process settings. Dynamical activity was originally proposed as an order parameter in the study of glass transitions [42]. The dynamical activity plays a central cost term in thermodynamic trade-off relations [18, 43, 20, 44]. For $K=1$, Eq. (12) reduces to the bound in Ref. [45]. While the right-hand side of Eq. (12) is written entirely in terms of quantities for a single process, the left-hand side contains the observable $N^{(\otimes K)}$ for the combined process of $K$ replicas. Note that Eq. (12) is an intermediate relation that will be used to obtain trade-off relations in the single process. In other words, although we consider the combined process of replicas $K$, this is purely for mathematical purposes, and the final results that will be shown later are relations for the single process [cf. Eqs. (16) and (17)]. This is the same idea as in methods such as multi-copy measurements and the replica trick. As shown in the introduction, in two-copy measurements, we use $\rho\otimes\rho$ but the quantity obtained from using these copies is $\mathrm{Tr}[\rho^{2}]$, which is a quantity referring to the single-copy case. Considering $r=1$ and $s\to\infty$ in Eq. (12), we obtain
@@
$\mathbb{E}\left[|N^{(\otimes K)}|\right]\leq N_{\max}^{(\otimes K)}\left(1-e^{-\tau K\mathfrak{a}(t=0)}\right),$ (14)
where $N_{\max}^{(\otimes K)}$ denotes the maximum value $N_{\max}^{(\otimes K)}\equiv\max_{\Gamma^{(\otimes K)}}\left|N^{(\otimes K)}(\Gamma^{(\otimes K)})\right|$. For $K=1$, the bound is identical to that reported in Ref. [45].

### v2

# Entropic Trade-off Relations in Stochastic Thermodynamics via Replica Markov Processes

@@ Abstract
Traditional thermodynamic trade-off relations usually apply to quantities that depend linearly on probability distributions. In contrast, many important information-theoretic measures, such as entropies, are nonlinear and thus difficult to analyze with existing frameworks. Motivated by replica ideas in quantum information and spin-glass theory, we introduce replica Markov processes, i.e., Markovian dynamics of $K$ independent, identical copies, and derive trade-off relations that bound relative moments of replica observables in terms of dynamical activity. By choosing appropriate replica observables, these inequalities translate into bounds on nonlinear quantities of the original single process. Focusing on entropic measures of uncertainty, we obtain upper bounds on the time derivative and the value of the Tsallis entropy for general trajectory-observable distributions, and corresponding bounds for diffusion on networks. We also illustrate how replicas can handle other nonlinear functionals such as extreme-value observables. Finally, we extend the construction to continuously monitored open quantum systems governed by Lindblad dynamics, where the bounds are expressed in terms of the quantum dynamical activity. These results provide entropic counterparts to activity-based uncertainty relations and establish a general method for constraining nonlinear information-theoretic quantities in stochastic and quantum thermodynamics.
I Introduction
Over the past decade, thermodynamic trade-off relations have been extensively studied in stochastic and quantum thermodynamics. These relations express a “no free lunch” principle in thermodynamics: faster, more precise operations require greater thermodynamic resources. In the thermodynamic uncertainty relations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], trajectory observables are considered and lower bounds on their relative variance are derived, where the bounds involve the entropy production and the dynamical activity. Extensions of these uncertainty relations to the quantum domain have been actively studied [17, 18, 21, 22, 23, 24]. Related work has also examined how to compute state-dependent expectation values [25, 26]. What these relations have in common is that the quantities being computed are linear in the underlying probabilities. For example, in thermodynamic uncertainty relations, we are often interested in quantities of the form
$\mathbb{E}[f]=\sum_{\Gamma}\mathcal{P}(\Gamma)\,f(\Gamma),$ (1)
where $\Gamma$ denotes a stochastic trajectory, $\mathcal{P}(\Gamma)$ is the trajectory probability, and $f(\Gamma)$ is an observable like a current. Equation (1) is linear with respect to probability $\mathcal{P}(\Gamma)$. Typical examples of Eq. (1) include dissipated heat and stochastic displacement. In contrast, nonlinear functions of the probabilities are generally much harder to compute, although they play a crucial role. A prominent example are generalized entropies such as Rényi and Tsallis, which measure the degree of uncertainty and the spread of information. The thermodynamic uncertainty relation expresses the uncertainty using the variance of an observable. However, in information theory, uncertainty is often quantified using entropy. When the output is categorical rather than numerical, the variance is not a natural measure of uncertainty. Entropy, by contrast, applies directly to categorical outcomes and can be used to quantify uncertainty in this setting. For example, when quantifying the extent that a random walker on a network has propagated over the network, variance is not suitable because the states are labels rather than numerical values. Establishing thermodynamic bounds for these nonlinear functionals remains a challenge because standard frameworks are inherently linked to linear expectations. There is thus a fundamental need for a theoretical bridge that maps nonlinear information measures onto the stochastic thermodynamics.
@@ I Introduction
Inspired by replica methods from quantum information and spin-glass theory, we introduce replica Markovian dynamics comprising $K$ identical copies of the process and derive bounds on quantities that depend nonlinearly on the underlying probability distributions. In quantum information and statistical mechanics, we often need to compute the expectation values of nonlinear physical quantities. For example, purity requires the square of the density operator, $\mathrm{Tr}[\rho^{2}]$. Although purity serves as a key quantity for distinguishing quantum behavior from classical behavior, its experimental measurement is known to be difficult. The swap trick [27] is a technique to measure purity without performing state tomography. The swap trick uses the following identity:
@@
where $\mathrm{SWAP}$ denotes the swap operator. The identity given in Eq. (2) shows that the purity $\mathrm{Tr}[\rho^{2}]$ can be obtained by preparing two replicas of $\rho$ and measuring the expectation value of the SWAP operator acting on $\rho\otimes\rho$. This avoids full state tomography of $\rho$, which would require many measurement shots to reliably estimate $\rho^{2}$. The use of replicas has been highly effective in classical shadow tomography [28, 29, 30] or in the study of the barren plateau in parametrized quantum circuits [31]. The application of hypothetical replicas to the evaluation of nonlinear quantities has a long history. In spin glasses [32, 33, 34], for example, we often need to evaluate the disorder average of the logarithm of the partition function $\mathbb{E}[\ln Z]$, which is difficult to compute directly. The replica method addresses this by introducing $n$ replicas of the system and expressing the logarithm as a limit $\ln Z=\lim_{n\to 0}(Z^{n}-1)/n$. This technique has been widely used in the analysis of learning theory [35, 36, 37, 38, 39] and conformal field theory [40, 41] . These examples illustrate the power of using replicas to calculate expectations of nonlinear quantities.
We introduce a method for deriving trade-off relations for nonlinear quantities of a single Markov process. This is achieved by considering $K$ independent identical copies (replicas) of the process and applying trade-off relations to suitably chosen replica observables. Through this approach, these nonlinear quantities are expressed as linear expectations within the replica framework. Crucially, replicas can be treated as purely virtual, analogous to the replica trick, allowing us to obtain fundamental constraints on the original single process. Our work then focuses on entropic measures of uncertainty. We derive upper bounds on both the time derivative and the value of the Tsallis entropy for general trajectory-observable distributions, relating these to dynamical activity [Eqs. (24) and (25)]. Additionally, we establish Tsallis bounds controlled by the initial activity [Eqs. (27) and (28)] Analogous results also obtained for network diffusion. Specifically, we obtain the time derivative [Eq. (31)] and the value of the Tsallis entropy for state distributions [Eq. (33)], where their bounds comprise the dynamical activity. Furthermore, we establish a bound on the Tsallis entropy of the state distribution, relating it to the initial dynamical activity [Eqs. (34) and (35)]. A notable finding is that these initial activity-controlled bounds depend solely on the local escape rate from the initial node [Eqs. (34) and (35)]. Beyond entropy, we demonstrate the utility of replicas for handling other nonlinear functionals, including extreme-value observables [Eqs. (38)–(41)]. Finally, we generalize this construction to continuously monitored open quantum systems governed by Lindblad dynamics, where the resulting bounds are formulated in terms of quantum dynamical activity [Eq. (54), (55) and (57)]. These results provide entropic counterparts to activity-based uncertainty relations. While conventional uncertainty relations typically lower-bound relative variances by the inverse activity, our bounds similarly show that entropy-based uncertainty is fundamentally constrained by dynamical activity.
@@
Figure 1: Illustration of Markov processes. (a) Single process scenario. This scenario is usually considered in stochastic thermodynamics. For a Markov process, we consider stochastic trajectories, random sample paths that represent individual realizations of the process. We are interested in stochastic quantities associated with the trajectories, whose statistics are linear in the trajectory probability. (b) $K=2$ replica process scenario. We consider two identical Markov processes. The state space is a product space of each Markov process and the trajectories are defined on the product space. We are interested in stochastic quantities of the replica process, which become nonlinear quantities in the original single process.
II Results
II.1 Replica Markov processes
@@
where the diagonal elements are defined by $W_{\mu\mu}=-\sum_{\mu^{\prime}\neq\mu}W_{\mu^{\prime}\mu}$. In stochastic thermodynamics, the trajectories generated by the master equation [Eq. (3)] are fundamental for analyzing the system’s stochastic dynamics. Several quantities, such as state displacement, dissipated heat, and Shannon entropy, are considered at the level of stochastic trajectories. We are interested in the dynamics within the time interval $0,\tau]$ ($\tau>0$). Suppose that there are $J$ jump events within the interval. Let $t_{j}$ be the time of the $j$-th jump and let $X_{j}\in\mathfrak{B}$ be the state after the $j$-th jump. The trajectory $\Gamma$ is then expressed by
@@
where $t_{0}=0$ is the initial time and $X_{0}$ is the initial state. Typically, we consider the counting observable:
$N(\Gamma)=\sum_{\nu,\mu(\nu\neq\mu)}C_{\nu\mu}N_{\nu\mu},$ (5)
where $N_{\nu\mu}$ is the number of jumps from $B_{\mu}$ to $B_{\nu}$ within the interval $0,\tau]$ and $C_{\nu\mu}$ are their associated real weights. As the first-level generalization, we can consider the observable $N_{\circ}(\Gamma)$. $N_{\circ}(\Gamma)$ can be an arbitrary function of $\Gamma$ as long as it satisfies the following condition:
$N_{\circ}(\Gamma_{\varnothing})=0,$ (6)
where $\Gamma_{\varnothing}$ represents a trajectory with no jump. Apparently, the counting observable in Eq. (5) satisfies the condition Eq. (6). Let $N_{\bullet}(\Gamma)$ be an arbitrary function defined on $\Gamma$. If we denote the sets of functions $N$, $N_{\circ}$, and $N_{\bullet}$ by $\{N\}$, $\{N_{\circ}\}$, and $\{N_{\bullet}\}$, respectively, then the inclusion holds:
$\{N\}\subset\{N_{\circ}\}\subset\{N_{\bullet}\}.$ (7)
As mentioned in the introduction, we consider replicas of the original process and regard the whole Markov process as a combined process (Fig. 1(b)). For simplicity, here we introduce the case of two replicas. That is, there are two identical processes, where there is no interaction between them. It is straightforward to extend this case to scenarios where there are $K$ replica processes. The state space is a product of the two state sets. The set of states in the combined Markov process of two replicas is given by $\mathfrak{B}^{(\otimes 2)}\equiv\{(B_{1}^{(1)},B_{1}^{(2)}),(B_{1}^{(1)},B_{2}^{(2)}),\ldots,(B_{D}^{(1)},B_{D}^{(2)})\}$. Therefore, the number of total states in the combined process is $|\mathfrak{B}^{(\otimes 2)}|=D^{2}$. From now on, when using a superscript $(k)$, it denotes a quantity associated with the $k$-th process, and when using a superscript $(\otimes K)$, it denotes a quantity for the combined system in which all processes $1,2,\dots,K$ are regarded as a single combined process. When we omit the superscript, the corresponding variables refer to those in the single process (i.e., $K=1$ case).
Let $p_{\mu\nu}^{(\otimes 2)}(t)$ be the probability that the combined system is in state $(B_{\mu}^{(1)},B_{\nu}^{(2)})\in\mathfrak{B}^{(\otimes 2)}$ at time $t$. Let $W_{\nu\mu}^{(k)}$ be the transition rate from state $B_{\mu}^{(k)}$ to state $B_{\nu}^{(k)}$ in the $k$-th process. Since we consider identical processes, $W_{\mu^{\prime}\mu}^{(k)}=W_{\mu^{\prime}\mu}$, we keep the superscript for clarity. The master equation corresponding to the combined process of the two replicas is given by
@@
II.2 Replica trade-off relations
Introducing the combined process with two replicas, we will focus on a general scenario in which there are $K$ replicas. We now introduce observables associated with trajectories. In the same way as in the single process case, we can consider observables $N^{(\otimes K)}(\Gamma^{(\otimes K)})$, $N_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})$, and $N_{\bullet}^{(\otimes K)}(\Gamma^{(\otimes K)})$. Specifically, $N^{(\otimes K)}$ is the counting observable, which yields the weighted sum of the number of jumps within the interval $0,\tau]$ in the $K$ replica process. The observable $N_{\circ}^{(\otimes K)}(\Gamma_{\varnothing}^{(\otimes K)})$ satisfies
$N_{\circ}^{(\otimes K)}(\Gamma_{\varnothing}^{(\otimes K)})=0,$ (13)
where $\Gamma_{\varnothing}^{(\otimes K)}$ is a trajectory in the combined process with no jump. The observable $N_{\bullet}^{(\otimes K)}$ is an arbitrary function of $\Gamma^{(\otimes K)}$.
Let $\mathbb{E}[X]$ denote the expectation value of a random variable $X$. Since $\Gamma^{(\otimes K)}$ is a random variable, observable $N^{(\otimes K)}(\Gamma^{(\otimes K)})$ is a random variable as well. From the uncertainty relation in Ref. [5], the following relation holds in the Markov process with $K$ replicas:
$\frac{\mathrm{Var}[N_{\bullet}^{(\otimes K)}]}{\tau^{2}\left(\partial_{\tau}\mathbb{E}[N_{\bullet}^{(\otimes K)}]\right)^{2}}\geq\frac{1}{K\mathcal{A}(\tau)}.$ (14)
Here, $\mathcal{A}(\tau)$ is the time-integrated dynamical activity:
$\mathcal{A}(\tau)=\int_{0}^{\tau}\mathfrak{a}(t)dt,$ (15)
where $\mathfrak{a}(t)$ is the dynamical activity at time $t$:
@@ II.2 Replica trade-off relations
Note that the dynamical activity $\mathfrak{a}(t)$ in Eq. (16) is a quantity of a single process [Eq. (3)]; the definition is the same as conventional Markov process settings. Dynamical activity was originally proposed as an order parameter in the study of glass transitions [42]. The dynamical activity plays a central cost term in thermodynamic trade-off relations [3, 43, 5, 44]. Since the replica process comprises $K$ independent processes, the dynamical activity of the combined process is $K\mathcal{A}(\tau)$, which directly yields Eq. (14).
Moreover, by employing another trade-off relation in Ref. [45], we obtain (see Appendix B for the derivation):
$\frac{\mathbb{E}\left[|N_{\circ}^{(\otimes K)}|^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[|N_{\circ}^{(\otimes K)}|^{r}\right]^{s/(s-r)}}\geq\frac{1}{1-e^{-\tau K\mathfrak{a}(t=0)}},$ (17)
where $0<r<s$. For $r=1$ and $s=2$, Eq. (17) reduces to the bound for the relative variance:
@@
$\frac{\mathrm{Var}\left[N_{\circ}^{(\otimes K)}\right]}{\mathbb{E}\left[N_{\circ}^{(\otimes K)}\right]^{2}}\geq\frac{1}{e^{K\tau\mathfrak{a}(t=0)}-1}.$ (18)
Equations (14) and (18) have similar forms, but there are several differences. First, the left-hand side of Eq. (14) is not written in the form of a relative variance, and the denominator on the left-hand side is given by a derivative of the expectation. By contrast, the left-hand side of Eq. (18) has the usual relative-variance form. In addition, the right-hand side of Eq. (14) is the reciprocal of the time-integrated dynamical activity $\mathcal{A}(\tau)$, while on the right-hand side of Eq. (18) the dynamical activity appears in the exponent, and only the dynamical activity $\mathfrak{a}(0)$, which is evaluated at $t=0$.
II.3 Entropic measures
While the right-hand sides of Eqs. (14) and (17) are written entirely in terms of quantities for a single process, the left-hand side contains the observable for the combined process of $K$ replicas. Note that Eqs. (14) and (17) are intermediate relations that will be used to obtain trade-off relations in the single process. In other words, although we consider the combined process of replicas $K$, this is purely for mathematical purposes, and the final results that will be shown later are relations for the single process. This is the same idea as in methods such as multi-copy measurements and the replica trick. As shown in the introduction, in two-copy measurements, we use $\rho\otimes\rho$ but the quantity obtained from using these copies is $\mathrm{Tr}[\rho^{2}]$, which is a quantity referring to the case of a single-copy.